# Periodic Korteweg-de Vries soliton potentials generate magnetic field strength with excellent quasisymmetry

### v3

@@ VI Numerical verification
Figure 4: The $B_{X}$ and $\iota$ profiles plotted for precise QA versions with negative shear (a) and positive shear (b).
Figure 5: Radially-summed two-term QS error, $f_{QS}$, plotted against average shear $\bar{\hat{s}}$ for different average rotational transforms, $\bar{\iota}$. The quasisymmetry error is summed over the $s=\{0.0,0.1,\dots 1.0\}$ surfaces. The definitions of the $\iota$ and $\hat{s}$ averages are given in (57). Note the V-shaped plots indicating a sharp change in $f_{QS}$ around a preferred $\bar{\hat{s}}$ for a particular $\bar{\iota}$, as predicted by our slab-geometry model. Note also the loss of preference for any particular $\bar{\hat{s}}$ as $\bar{\iota}$ becomes smaller.
@@
Figure 6: $(\partial_{l}B)^{2}$ as a function of $B$, on two surfaces, across the four stages of optimization for the Landreman-Paul precise QA configuration.
@@
Figure 13: $B$ on a magnetic surface parameterised by $l$ and $t$
Figure 14: $(\partial_{l}B)^{2}$ as a function of $B$ on two surfaces, on configurations at three different mean rotational transforms. Only at low rotational transform and far away from the axis, as in d), is a cubic a poor fit.
@@
We shall now demonstrate numerically that a large class of quasisymmetric $B$ satisfies the integrated form of KdV as given in (52). As demonstrated in Fig. 1(a), $(dB/dl)^{2}$, when plotted as a function of $B$, has the form of a cubic polynomial in the precise QA equilibrium Landreman and Paul (2022); the scatter is due to imperfect QS. On the other hand, Fig. 1(b) shows that increasing the polynomial order beyond three does not lead to an observable decrease in error (defined as $1-r^{2}$, where $r^{2}$ is the coefficient of determination), as the fit error is dominated by QS error. Thus, Fig. 1 supports equation (52) as a model for a device with perfect QS. A similar analysis for other configurations, which tend to have more QS error, is shown in Appendix E. (Note that $s$ is the normalized toroidal flux in all figures.)
In Fig. 2, we demonstrate that $B$ appears roughly sinusoidal far from the separatrix but becomes soliton-like (sharper peaks, flatter troughs) as the separatrix is approached. Note that the $B$ shown in Fig. 2 is perfectly quasisymmetric, as symmetry-breaking modes have been filtered out. The period of $B$ increases as $L(\psi)\to\infty$ on the separatrix.
With the global description of the exact quasisymmetric $B$, (55), at our disposal, we now investigate a key element of this theory: the physical meaning of the roots $B_{i}$. In Fig. 3(b), we show the typical behavior of the roots. Symmetry-breaking modes have been filtered out, and the roots are obtained by a cubic fit of the numerical data. Points represent the actual minimum and maximum values of $B$ on each flux surface. Surprisingly, the NAE description of $B_{M}$ and $B_{m}$ continues to hold even far from the axis. As shown in Fig. 3(a), the divergence in the connection length as the two roots meet is indeed due to the formation of an X-point that leads to a chain of islands.
The third root $B_{X}$ depends quite sensitively on the $\iota$ profiles (Fig. 4). We show in Fig. 4 that for small but finite shear, $B_{X}$ tracks the $\iota$ profile quite closely but varies more for positive shear. This is in accordance with our free-surface approach to QS, which showed that $B_{X}$ must be a function of the shear parameter.
The question of why quasisymmetric devices with excellent QS, such as the Landreman-Paul solution, have a small but finite magnetic shear is important to understand both for theoretical and practical reasons since shear controls MHD stability. However, NAE for QS is not very helpful when it comes to shear, as we discussed earlier. We shall now present an explanation based on the insight from the slab geometry model and through optimization with target shears. In Fig. 5, we show how the quality of QS $f_{QS}$ depends on the average shear for a given average rotational transform. The average shear is computed by fitting a line to the $\iota$ profile
$\iota(s)\approx\iota_{0}+\iota_{1}s,\quad\bar{\hat{s}}\equiv-\iota_{1}/\bar{\iota},$ (57)
where $\bar{\iota}$ is the average of the actual $\iota$ profile. Since these QS devices have very low shear, this definition suffices. From an optimization point of view, asking for a specific $\bar{\hat{s}}$ for a specific average $\iota$ will generally lead to degradation in the QS quality because of the extra shear constraint. However, from the sharpness of the V-shaped plots in Fig. 5, we observe that QS seems to have a strong preference for a particular $\bar{\hat{s}}$ for a given $\bar{\iota}$, particularly in the range of $\bar{\iota}\geq 0.32$. For lower values of $\bar{\iota}$, the V-shape flattens out. In light of the slab-geometry model, we can explain this feature as a direct consequence of the overdetermination problem. The force balance and the QS condition can only be satisfied simultaneously for a narrow range of shear around a preferred value of $\iota$ that is not too small.
Next, we demonstrate that the maximum volume that a quasisymmetric field can occupy can be estimated by the crossings of the roots of $\partial B/\partial l$ (52). From the analytical solution for $B$ (55), we find that the periodicity of $B$ is maintained as long as $10\leq m<$ 1. As we saw earlier, the limit $(m\to 1,B_{m}\to B_{X})$ was the limit in which the connection length diverges. Using the simple NAE estimates for the roots $B_{m}$ and $B_{X}$, we can estimate where the two roots will intersect, thereby signaling the breakdown of periodicity. We have shown in Fig. 3 how this estimate can predict the maximum volume of nested surfaces in precise QA extended using coils developed recently Wechsung et al. (2022). We note that the appearance of the resonant 4/10 island chain in Fig. 3 is not due to a resonance with the helicity of the QA Rodríguez and Bhattacharjee (2021). We further note that since the non-QS modes were filtered out, the breakdown of the topology of the field strength does not play any role in the divergence of the connection length.
Lastly, we show how the integrated form of KdV is approached through the stages of optimization, where the parameter space is progressively expanded. In Fig. 6, for the Landreman-Paul precise QA configuration, not only does the ‘noise’ decrease (as expected from improving QS), but the shape gets increasingly cubic. We show this both for a surface close to the axis and a surface close to the edge. Figures 7 and 8 further show how the KdV integrated cubic increases in accuracy (as determined by the coefficient of determination) further into the optimization process. In Figures 9, 10, 11, and 12, we show similar plots for the low mean rotational transform configurations introduced in Buller et al. (2024). Note that the trend is towards quartic instead of cubic. As evident from these plots, the PP is established in the entire volume within the very first few stages of optimization and is maintained thereafter.
VII Data driven verification
@@
We have put forth several analytical arguments in favor of the connection between QS and the well-known integrable PDEs, KdV, and Gardner’s equation (for small rotational transforms). Our analysis so far has been based on NAEs, an approximate slab geometry model calculation and its extensions to more realistic geometries, and finally, soliton theory. We have also demonstrated numerically good agreement with several configurations with good QS. In this Section, we take a different approach based on data-driven techniques to determine the underlying system of equations describing quasisymmetric $B$ on a flux surface. Our primary tool is pySINDy de Silva et al. (2020); Kaptanoglu et al. (2022), which fits for governing differential equations and favors fewer terms to combat over-fitting.
@@
We have primarily used the family of quasi-axisymmetric VMEC equilibria generated in Buller et al. (2024). For additional verification, we considered further configurations from the QUASR database Giuliani (2023). Henceforth, we shall refer to them as the Buller and the Giuliani configurations. For the latter, configurations were selected with three representative mean rotational transforms. They were chosen to have the number of field periods, $N_{fp}=2$, an aspect ratio close to six (to mirror the Buller configurations), and the lowest possible quasi-axisymmetry error (as defined in Giuliani (2023)).
[…]
@@
Next, we use pySINDy to discover the underlying ODEs that govern $B$ on a flux surface. To that end, a sequentially thresholded least squares algorithm Brunton et al. (2016); Zhang and Schaeffer (2019) was used for the fitting, with a cost function that promotes sparsity in the data fit. For each magnetic configuration’s flux surface, the $B$ against $l$ profiles of multiple field lines spanning the surface were used as training data. The quantity $(\partial_{\ell}B)^{2}$ on a flux surface was fit against a linear combination of terms from a ‘feature library,’ chosen here to be a selection of polynomial terms, with the model parameters as the coefficients. Since the KdV and Gardner’s equation, together with the traveling waveform, lead to cubic and quartic polynomials, using the polynomial basis in pySINDy was a natural choice. A range of models was produced at varying sparsity parameters; the models were then tested on data from magnetic field lines not used in the training data. For sparsity parameters that were too large, it was found that there was a sharp drop in the success of the fit as too many terms were dropped; we chose the pySINDy model before this drop in performance. This is a version of hyperparameter scanning to find the Pareto-optimal fit (optimal tradeoff of model complexity versus model accuracy) to the data.
@@
In Fig. 14, we demonstrate the outcome of the pySINDy fit to $(\partial_{\ell}B)^{2}$ on a flux surface for three representative configurations. A cubic was a good fit for all the configurations on flux surfaces sufficiently close to the axis, as expected from the NAE analysis. We see in Fig. 15 that the shapes from the Giuliani configurations are similar to those of the Buller configurations.
@@
For configurations with low mean rotational transform, a cubic was no longer the best model on the outermost surfaces (see Figures 16 and 17); pySINDy indicated that a quartic was an effective fit on such surfaces. We also see that a quintic performs well. However, within the error range, the benefit of using a quintic is very marginal compared to the quartic, as demonstrated clearly in Figs. 16 and 17. In Fig. 18, we see that the deviation from the mean from many pySINDy models generated by sub-sampling is small. This provides additional evidence that cubic models fit the data extremely well, except near the plasma boundary in low-iota configurations, for which a quartic correction is robustly picked up during model identification.
@@
Finally, the evolution of the roots in the representative sample configurations and surfaces is shown in Fig. 19. For the cubic case described by KdV, there are three real roots, as we discussed earlier in Section VI. The case for low $\iota$, described by Gardner’s equation, on the other hand, can have two real and two complex conjugate roots owing to its quartic nature. In the leftmost panel of Fig. 19, we show the two real roots of the quartic $(\partial_{\ell}B)^{2}$ given by the maximum and the minimum of $B$ on a flux surface $s$ and the other pair of complex conjugate roots.
VIII Discussions
In summary, we have demonstrated that periodic soliton solutions of the KdV equation lead to a broad class of excellent quasisymmetric $B$. We have done so by employing three different approaches. In the first approach, we reformulate the QS problem as a free-surface problem. The inner (closer to the axis) solution is assumed to be the second-order NAE solution, which has been demonstrated to capture essential features of even low-aspect-ratio practical designs. The outer free-surface solution is then constructed by considering the length scale over which the magnetic shear varies. We can perturbatively solve the overdetermination problem by ensuring compatibility of the force balance and the QS condition. Our results show that shear is constrained to have an optimum value by the overdetermination problem and the magnetic field strength. The field strength must also satisfy the KdV equation, a well-known integrable PDE. We can also obtain Gardner’s equation for small values of the rotational transform.
@@ VIII Discussions
In the second approach, we show that under three very general assumptions, namely, periodicity in $\ell$, analyticity of $B$, and the existence of a nontrivial TW frame, we can motivate why QS naturally leads to the KdV or Gardner’s equation for the magnetic field strength. The essential property is the Painlevé property, which ensures $B$ stays single-valued. Several results follow from the connection between QS and the KdV equation.
@@
Firstly, the description of a class of precisely quasisymmetric $B$ can be reduced to understanding the behavior of the roots $B_{i}$ of a cubic polynomial. These roots correspond to extrema or saddle points of $B$ on a surface. In particular, $B$, which satisfies the periodic KdV equation, is the potential of the periodic Schrodinger equation where $B_{i}$’s are related to the endpoints of the band gap (Kamchatnov, 2000; Novikov et al., 1984). Since the spectrum of KdV is time-independent, the roots do not depend on time. This confirms the well-known statement that the extrema or saddle points of $B$ depend only on the flux label in QS Helander (2014). Moreover, the free-surface approach showed that $B_{X}$ is a function of the shear. Therefore, the invariance of $B_{X}$ on a flux surface effectively implies that the rotational transform $\iota(\psi)$ must be independent of the field line label (time). In other words, QS leads to an integrable field line Hamiltonian, which preserves nested flux surfaces (Helander, 2014). Furthermore, since various aspects of the flux surface shaping can be deduced (Rodriguez, 2022) from the Fourier coefficients of $B$, the hidden lower dimensionality is beneficial for understanding shaping. A similar hidden lower dimensionality is found in the description of the curvature of isodynamic magnetic field lines (Schief, 2003). We repeat for mphasis that these hidden lower dimensionalities in the field strength or geometry are a direct consequence of the underlying soliton dynamics: the KdV equation in the case of QS and vortex filament equation in the case of isodynamic fields.
@@
Thirdly, we can estimate the maximum toroidal volume QS allows solely from the properties of the roots and independent of any particular geometry. The overlap of the third root with the second signals the breakdown of periodicity of $B$, as demonstrated numerically. At this radius, a separatrix could form, which could potentially be used as the basis for a non-resonant divertor (Boozer, 2015; Punjabi and Boozer, 2020).
@@
Our third approach is data-driven. Utilizing the vast amount of data on quasisymmetric equilibria generated in Buller et al. (2024) and the QUASR database Giuliani (2023), and the state-of-the-art tool pySINDy, which uses sparse regression techniques to discover underlying models, we recover KdV and Gardner’s equation. Thus, the Painlevé property that demands that $(\partial_{\ell}B)^{2}$ can be at most a quartic in $B$ is verified accurately. This data-driven approach has the advantage that it can be applied to any configuration (including when quasi-symmetry is not relevant) to find low-dimensional ODE or PDE models underlying the data.
Given the well-known robust stability of soliton solutions Tao (2009), the spectral stability of the cnoidal waves (Bottman and Deconinck, 2009), and the robustness of the single-valuedness criterion, which led to the Painlevé property, perhaps explain why any numerical optimizer trying to find good QS using the 2-term form or the triple product form will most likely find only cubic or quartic terms in $(\partial_{\ell}B)^{2}$.

### v4

@@ V Numerical verification
Figure 5: The $B_{X}$ and $\iota$ profiles plotted for precise QA versions with negative shear (a) and positive shear (b). Symmetry-breaking modes have been filtered out.
@@
Figure 6: $(\partial_{l}B)^{2}$ as a function of $B$, on two surfaces, across the four stages of optimization for the Landreman-Paul precise QA configuration. Top row: surface near edge at $s=0.96$. Bottom row: surface near core at $s=0.1$. Left to right: increasing stages in the Fourier continuation based optimization Landreman and Paul (2022).
@@
We now demonstrate numerically that a large class of quasisymmetric $B$ satisfies the integrated form of KdV as given in (22). As demonstrated in Fig. 2(a), $(dB/dl)^{2}$, when plotted as a function of $B$, has the form of a cubic polynomial for the precise QA equilibrium Landreman and Paul (2022); the scatter is due to imperfect QS. On the other hand, Fig. 2(b) shows that increasing the polynomial order beyond three does not lead to an observable decrease in error (defined as $1-r^{2}$, where $r^{2}$ is the coefficient of determination), as the fit error is dominated by QS error. Thus, Fig. 2 supports equation (22) as a model for a device with perfect QS. A similar analysis for other configurations, which tend to have greater QS errors, is shown in Appendix C. (Note that $s$ is the normalized toroidal flux in all figures.)
With the global description of the exact quasisymmetric $B$, (34), at our disposal, we now investigate a key element of the theory: the physical meaning of the roots $B_{i}$. In Fig. 4(b), we show the radial dependence of the roots. Symmetry-breaking modes have been filtered out, and the roots are obtained by a cubic fit of the numerical data. Points represent the actual minimum and maximum values of $B$ on each flux surface. Surprisingly, the NAE description of $B_{M}$ and $B_{m}$ continues to hold even far from the axis. As shown in Fig. 4(a), the location where $B_{m}$ and $B_{X}$ meet coincides with the location of an X-point that leads to a chain of islands. This is consistent with periodicity breaking down for $m\to 1$.
The third root $B_{X}$ depends quite sensitively on the $\iota$ profiles (Fig. 5). We show in Fig. 5 that, for small but finite shear, $B_{X}$ tracks the $\iota$ profile quite closely but varies more for positive shear.
Lastly, we show how the integrated form of KdV is approached through the stages of optimization, where the parameter space is progressively expanded. In Fig. 6, for the Landreman-Paul precise QA configuration, not only does the ‘noise’ decrease (as expected from improving QS), but the shape becomes increasingly cubic. We show this both for a surface close to the axis and a surface close to the edge. Figures 7 and 8 further show how the integrated KdV cubic increases in accuracy (as determined by the coefficient of determination) further into the optimization process. In Figures 9, 10, 11, and 12, we show similar plots for the low mean rotational transform configurations introduced in Buller et al. (2025). Note that the trend is towards quartic instead of cubic. As evident from these plots, the PP is established in the entire volume within the very first few stages of optimization and is maintained thereafter.
VI Data driven verification
Figure 13: $B$ on a magnetic surface parameterised by $l$ and $t$, for precise-QA-like configurations with varying rotational transformBuller et al. (2025). Top row: flux surface close to the core ($s=0.1735$. Bottom row: flux surface close to the edge $s=0.9898$.
Figure 14: $(\partial_{l}B)^{2}$ as a function of $B$ on two surfaces, for precise-QA-like configurations with varying rotational transformBuller et al. (2025). Only at low rotational transform and far away from the axis, as in d), is a cubic a poor fit.
@@ VI Data driven verification
We have put forth several analytical arguments in favor of the connection between QS and the well-known integrable PDEs, KdV, and Gardner’s equation (for small rotational transforms). We have also demonstrated numerically good agreement with several configurations with good QS. In this Section, we take a different approach based on data-driven techniques to determine the underlying system of equations describing quasisymmetric $B$ on a flux surface. Our primary tool is pySINDy de Silva et al. (2020); Kaptanoglu et al. (2022), which fits governing differential equations and favors fewer terms to combat over-fitting.
@@
We have primarily used the family of quasi-axisymmetric VMEC equilibria generated in Ref. Buller et al. (2025). For additional verification, we considered further configurations from the QUASR database Giuliani (2023). Henceforth, we refer to them as the Buller and the Giuliani configurations. For the latter, configurations were selected with three representative mean rotational transforms. They were chosen to have the number of field periods, $N_{fp}=2$, an aspect ratio close to six (to mirror the Buller configurations), and the lowest possible quasi-axisymmetry error (as defined in Giuliani (2023)).
Using the Buller configurations, we verified the quasisymmetry-imposed traveling wave form (12). We calculated $H=-\partial_{\alpha}B/\partial_{l}B$, taking $l=0$ at $\phi=0$ and considering one field period. The traveling waveform could then be verified by the fact that $H$ was independent of $l$. Equivalently, we calculated $t=\int_{0}^{\alpha}d\alpha^{\prime}H$ and plotted $B$ on a surface parameterized by $l$ and $t$, in which case equation (12) indicates that we should see traveling waves of unit speed. These results are presented in Fig. 13, where we can clearly see QS manifest as a traveling wave. With the traveling waveform, the PDE governing $B$ reduces to ODEs, which pySINDy can easily determine. In fact, we also attempted to use pySINDy to identify PDEs directly, and could straightforwardly recover the traveling wave. Even in configurations with relatively substantial deviations from quasi-symmetry, the traveling wave structure is quite robust. Under the traveling wave assumption, the equations are reduced to ODEs anyway, so we focused on the ODE discovery problem. Future work can try to use pySINDy for PDE identification in regimes where significant deviations from traveling wave behavior are observed, which might require configurations quite far from quasi-symmetry.
@@
Next, we use pySINDy to discover the underlying ODEs that govern $B$ on a flux surface. To that end, a sequentially thresholded least squares algorithm Brunton et al. (2016); Zhang and Schaeffer (2019) was used for the fitting, with a cost function that promotes sparsity in the data fit. For each magnetic configuration’s flux surface, the $B$ against $l$ profiles of multiple field lines spanning the surface were used as training data. The quantity $(\partial_{\ell}B)^{2}$ on a flux surface was fitted against a linear combination of terms from a ‘feature library,’ chosen here to be a selection of polynomial terms, with the model parameters as the coefficients. Since the KdV and Gardner’s equation, together with the traveling waveform, lead to cubic and quartic polynomials, using the polynomial basis in pySINDy was a natural choice. A range of models was produced at varying sparsity parameters; the models were then tested on data from magnetic field lines not used in the training data. For sparsity parameters that were too large, it was found that there was a sharp drop in the success of the fit as too many terms were dropped; we chose the pySINDy model before this drop in performance. This is a version of hyperparameter scanning to find the Pareto-optimal fit (optimal tradeoff of model complexity versus model accuracy) to the data.
@@
In Fig. 14, we demonstrate the outcome of the pySINDy fit to $(\partial_{\ell}B)^{2}$ on a flux surface for three representative configurations. A cubic was a good fit for all the configurations on flux surfaces sufficiently close to the axis, as expected from the NAE analysis.
@@
For configurations with low mean rotational transform, a cubic was no longer the best model on the outermost surfaces (see Figures 16 and 17); pySINDy indicated that a quartic was an effective fit on such surfaces. We also see that a quintic performs well. However, within the error range, the benefit of using a quintic is very marginal compared to the quartic. We see in Fig. 15 that the Giuliani configurations display similar behavior with deviation from a cubic fit away from the axis for configurations with low mean-$\iota$. In Fig. 18, we see that the deviation from the mean from many pySINDy models generated by sub-sampling is small. This provides additional evidence that cubic models fit the data extremely well, except near the plasma boundary in low-iota configurations, for which a quartic correction is robustly picked up during model identification.
@@
Finally, the evolution of the roots in the representative sample configurations and surfaces is shown in Fig. 19. For the cubic case described by KdV, there are three real roots, as we discussed earlier in Section V. The case for low $\iota$, described by Gardner’s equation, on the other hand, can have two real and two complex conjugate roots owing to its quartic nature. In the leftmost panel of Fig. 19, we show the two real roots of the quartic $(\partial_{\ell}B)^{2}$ given by the maximum and the minimum of $B$ on a flux surface $s$ and the other pair of complex conjugate roots.
VII Discussions
@@
In summary, we have demonstrated that periodic soliton solutions of the KdV equation lead to a broad class of excellent quasisymmetric $B$. We show that under three very general assumptions, namely, periodicity in $\ell$, analyticity of $B$, and the existence of a nontrivial TW frame, we can motivate why QS naturally leads to the KdV or Gardner’s equation for the magnetic field strength. The essential property is the Painlevé property, which ensures $B$ stays single-valued. Several results follow from the connection between QS and the KdV equation.
@@
Firstly, the description of a class of precisely quasisymmetric $B$ can be reduced to understanding the behavior of the roots $B_{i}$ of a cubic polynomial. These roots correspond to extrema or saddle points of $B$ on a surface. In particular, $B$, which satisfies the periodic KdV equation, is the potential of the periodic Schrodinger equation, where $B_{i}$’s are related to the endpoints of the band gap (Kamchatnov, 2000; Novikov et al., 1984). Since the spectrum of KdV is time-independent, the roots do not depend on time. This confirms the well-known statement that the extrema or saddle points of $B$ depend only on the flux label in QS Helander (2014). Therefore, the invariance of $B_{X}$ on a flux surface effectively implies that the rotational transform $\iota(\psi)$ must be independent of the field line label (time). In other words, QS leads to an integrable field line Hamiltonian, which preserves nested flux surfaces (Helander, 2014). Furthermore, since various aspects of the flux surface shaping can be deduced (Rodriguez, 2022) from the Fourier coefficients of $B$, the hidden lower dimensionality is beneficial for understanding shaping. A similar hidden lower dimensionality is found in the description of the curvature of isodynamic magnetic field lines (Schief, 2003). We repeat for emphasis that these hidden lower dimensionalities in the field strength or geometry are a direct consequence of the underlying soliton dynamics: the KdV equation in the case of QS and the vortex filament equation in the case of isodynamic fields.
@@
Thirdly, we can estimate the maximum toroidal volume that QS allows solely from the properties of the roots and independent of any particular geometry. The overlap of the third root with the second could signal the breakdown of periodicity of $B$, as demonstrated numerically. At this radius, a separatrix could form, which could potentially be used as the basis for a non-resonant divertor (Boozer, 2015; Punjabi and Boozer, 2020).
@@
Our final approach is data-driven. Utilizing the vast amount of data on quasisymmetric equilibria generated in Buller et al. (2025) and the QUASR database Giuliani (2023), and the state-of-the-art tool pySINDy, which uses sparse regression techniques to discover underlying models, we recover KdV and Gardner’s equation. Thus, the Painlevé property that demands that $(\partial_{\ell}B)^{2}$ can be at most a quartic in $B$ is verified accurately. This data-driven approach has the advantage that it can be applied to any configuration (including when quasi-symmetry is not relevant) to find low-dimensional ODE or PDE models underlying the data.
Given the well-known robust stability of soliton solutions Tao (2009), the spectral stability of the cnoidal waves (Bottman and Deconinck, 2009), and the robustness of the single-valuedness criterion, which led to the Painlevé property, this perhaps explains why any numerical optimizer trying to find good QS using the 2-term form or the triple-product form will most likely find only cubic or quartic terms in $(\partial_{\ell}B)^{2}$.